OpenStudy (anonymous):

Y = 2x + 4 is it a non-proportional or proportional linear relationship? how do i solve this?

OpenStudy (anonymous):

plz help me

OpenStudy (unklerhaukus):

if y is proportional to x \[y\propto x\\y=kx\] where \(x\) is a constant

OpenStudy (anonymous):

i dont know

OpenStudy (anonymous):

I believe its non proportional if you are solving for x then it is x= (y-4)/2 or simplifying then its y= 2(x+2)

OpenStudy (unklerhaukus):

the plus four term makes the expression non-proportional

OpenStudy (anonymous):

y?

OpenStudy (unklerhaukus):

because for a proportional relationship the line must pass through the origin ie y(0) must equal 0

OpenStudy (anonymous):

like 2 times 0 equals 0 but +4 is 4 so its non proportinal

OpenStudy (unklerhaukus):

yeah